Two locomotives are 150 meters apart each having a mass of 4.0x10^5 kg. Find the gravitational force that exist between them.


1
Expert's answer
2021-12-12T16:40:17-0500

The force can be found by the law of universal gravitation:


"F=GM·M/R^2=GM^2/R^2=4.75·10^{-4}\text{ N}."